OpenStudy (anonymous):

What is the solution to the equation 7 ^–3x ≈ 9 I have no idea where to start please help

OpenStudy (ivancsc1996):

First natural log:\[\ln (7^{-3x})=-3x \ln (7)\approx 9 \rightarrow x=-\frac{ 9 }{ 3\ln (7) }=-\frac{ 3 }{ \ln (7) }\]

OpenStudy (anonymous):

Okay